# Strong Decays of the Light Exotic $0^{+-}$ and $2^{+-}$ Hybrid Mesons

### v2

@@ III.3 Isoscalars
The light $0^{+-}$ isoscalar hybrid appears to be broad, with a very large contribution from the $b_{1}(1235)\pi$ channel and smaller, but significant, contributions from $h_{1}(1170)\eta$ and $K_{1L}\bar{K}$, and it might therefore be hard to detect. On the other hand, the strange $0^{+-}$ is narrow, and could appear as a sharp peak in the $K_{1L}\bar{K}$ channel. The $K(1460)\bar{K}$ channel is below threshold for the light state and negligible for the strange state. Again, the reason for this lies in the small value of $\beta_{m}$ parameter obtained for the $2^{1}S_{0}$ $K(1460)$ state, just as it was the case for the $\pi(1300)$ above, which appears to be something unique to our model. In fact, both Page:1998gz and Chen2025 report very large contributions coming from $K(1460)\bar{K}$, with the latter reporting a branching ratio of $\frac{\Gamma(K(1460)\bar{K})}{\Gamma_{tot}}=\frac{394.4}{484}=81.5\%$.
@@
The strange $2^{+-^{\prime}}$ is the narrowest state, with $K_{1H}\bar{K}$ being the only significant decay mode. Our predictions here are more in line with the PSS model, while Chen2025 reports larger widths.
@@ IV Summary and Conclusions
We have developed a model of hybrid structure and light hybrid decays that leverages the QCD Hamiltonian in Coulomb gauge and assumes a single constituent gluon as a good description of gluodynamics. This approach finds support in lattice field computations, which determine that an axial gluonic degree of freedom is sufficient to describe the main features of the low lying hybrid spectrum. The results presented here are compared to previous work, such as the flux-tube PSS model Page:1998gz and a recent constituent gluon model by Chen and Liu Chen2025 which is in many regards very similar to our approach. We find that our results mostly agree with those papers, with some exceptions. The most notable difference is the $0^{+-}$ which is very narrow in our computation, due to the suppression of the decay mode $0^{+-}\rightarrow 2^{1}S_{0}+1^{1}S_{0}$. The works cited above find that particular decay mode to make the largest contribution to the total width, both for the isovector and the strange isoscalar.
@@
Additionally, the total widths are clearly dependent on the hybrid masses used. The masses were crudely extrapolated from the lattice computations by shifting all of them down by the same amount so that the $1^{-+}$ mass agrees with the mass of the $\pi(1600)$, which was set at 1600 MeV. Choosing different masses would clearly change the total widths values and could allow additional channels to go above or below threshold. This is a very important point in assessing the reliability of the results, because, as obtained by lattice computations, all the states of interest are actually decaying resonances and the masses determined should all be considered to have a systematic error at least as big as the (unknown) hadronic width of the states jjd .
@@
Lastly, all of the strange isoscalar hybrids appear very narrow. This is in part because we have chosen ideal light-strange mixing, which only allows kaon-pair decay modes.
@@
To our knowledge, there is currently no experimental evidence of $0^{+-}$ exotic states, but preliminary results for the reaction $\pi^{-}p\rightarrow 2\pi^{-}2\pi^{+}\pi^{0}$ by the BNL(E852) collaboration, found a strong 1.9 GeV signal for a $I^{G}J^{PC}=0^{-}2^{+-}$ state (labeled $h_{2}$) in the process $\pi^{-}p\rightarrow h_{2}n\rightarrow b_{1}\pi n\rightarrow 2\pi\omega\rightarrow 5\pi n$ Adams:2005tx . This could be a candidate for our isovector $2^{+-}$ state, whose mass is very close to 1.9 GeV and whose main decay mode is the $b_{1}(1235)\pi$. However, these preliminary results need confirmation and should therefore be taken with caution. Searches at GlueX and in multi-pion final states in $J/\psi$ and $\chi_{c1}$ decay at BES appear to be promising for providing such confirmation.
@@
The most obvious continuation of the present work would be to compute the strong decays of the remaining states in Table 2. These hybrid states all have conventional quantum numbers ($(0,2)^{++},(1,3)^{+-}$) and can therefore mix with conventional mesons, making their identification more challenging. Hence, computing their strong decays might reveal some interesting features that could distinguish them from conventional mesons and thus aid in their detection.
Additionally, the model presented here can be extended to include spin-dependent terms, which in turn could be used to compute the mass spectrum of hybrids.

### v3

@@ III.3 Isoscalars
The light $0^{+-}$ isoscalar hybrid appears to be broad, with a very large contribution from the $b_{1}(1235)\pi$ channel and smaller, but significant, contributions from $h_{1}(1170)\eta$ and $K_{1L}\bar{K}$, and it might therefore be hard to detect. On the other hand, the strange $0^{+-}$ is narrow, and could appear as a sharp peak in the $K_{1L}\bar{K}$ channel. The $K(1460)\bar{K}$ channel is below threshold for the light state and negligible for the strange state. Again, the reason for this lies in the small value of the $\beta_{m}$ parameter obtained for the $2^{1}S_{0}$ $K(1460)$ state, just as it was the case for the $\pi(1300)$ above, which appears to be something unique to our model. In fact, both Page:1998gz and Chen2025 report very large contributions coming from $K(1460)\bar{K}$, with the latter reporting a branching ratio of $\frac{\Gamma(K(1460)\bar{K})}{\Gamma_{tot}}=\frac{394.4}{484}=81.5\%$.
@@
The strange $2^{+-^{\prime}}$ is the narrowest state, with $K_{1H}\bar{K}$ being the only significant decay mode. Our predictions here are more in line with the PSS model, while Ref. Chen2025 reports larger widths.
@@ IV Summary and Conclusions
We have developed a model of hybrid structure and light hybrid decays that leverages the QCD Hamiltonian in Coulomb gauge and models the relevant gluodynamics in terms of a dynamically generated constituent gluon. This approach finds support in lattice field computations, which determine that an axial gluonic degree of freedom is sufficient to describe the main features of the low lying hybrid spectrum. The results presented here are compared to previous work, such as the flux-tube PSS model Page:1998gz and a recent constituent gluon model by Chen and Liu Chen2025 which is in many regards very similar to our approach. We find that our results mostly agree with those papers, with some exceptions. The most notable difference is the $0^{+-}$ which is very narrow in our computation, due to the suppression of the decay mode $0^{+-}\rightarrow 2^{1}S_{0}+1^{1}S_{0}$. The works cited above find that particular decay mode to make the largest contribution to the total width, both for the isovector and the strange isoscalar.
@@
Additionally, the total widths are clearly dependent on the hybrid masses used. The masses were crudely extrapolated from the lattice computations by shifting all of them down by the same amount so that the $1^{-+}$ mass agrees with the mass of the $\pi(1600)$, which was set at 1600 MeV. Choosing different masses would clearly change the total widths and could allow additional channels to go above or below threshold. This is a very important point in assessing the reliability of the results, because, as obtained by lattice computations, all the states of interest are actually decaying resonances and the masses determined should all be considered to have a systematic error at least as big as the (unknown) hadronic width of the states jjd .
In order to do a sensitivity analysis of our results, we computed the width variation as a function of several of the parameters that are used in the model for the isovector $0^{+-}\rightarrow\pi\pi(1300)$, the $H_{2}$ isoscalar $2^{+-}\rightarrow b_{1}(1235)\pi$, and for the total width of the $H_{4}$ isovector $2^{+-}$. The parameters varied were the Coulombic strength $\alpha_{S}$, the hybrid mass $M_{hyb}$, the hybrid variational parameters $\alpha_{hyb}$ and $\beta_{hyb}$, and the meson parameters $\beta_{m}$. In each case, we varied each parameter by $\pm 10\%$, but for the total $2^{+-}$ width, all the meson betas were set to be equal, ranging from $0.30$ to $0.50$ GeV. The results are shown in Figures 3, 4, and 5. As the plots show, the decays are most sensitive to the $\beta$ parameters and the hybrid masses. The dependence on $\alpha_{S}$ is simple – $\Gamma$ is largely proportional to $\alpha_{S}$ – but less significant. These results are expected in large part. A higher hybrid mass increases the width value by increasing the available phase space for the decay and also by allowing additional channels to open up. A variation in the beta parameters for both the hybrids and conventional mesons, affects the widths by changing the value of the Gaussian functions in the radial expansion, which depending on the $\beta$ value, can have a large effect on the width.
Figure 3: Width variation of $0^{+-}\rightarrow\pi\pi(1300)$ as a function of model parameters. Parameters were varied by 10%.
Figure 4: Width variation of the $H_{2}$ isoscalar $2^{+-}\rightarrow b_{1}(1235)\pi$ as a function of model parameters. Parameters were varied by 10%.
Figure 5: Total width variation of the $H_{4}$ isovector $2^{+-}$ as a function of model parameters. Parameters were varied by 10%, except for the $\beta_{mes.}$ which were varied between 0.30 GeV and 0.50 GeV.
@@
We note that all of the strange isoscalar hybrids appear very narrow. This is in part because we have chosen ideal light-strange mixing, which only allows kaon-pair decay modes.
@@
To our knowledge, there is currently no experimental evidence of $0^{+-}$ exotic states, but preliminary results for the reaction $\pi^{-}p\rightarrow 2\pi^{-}2\pi^{+}\pi^{0}$ by the BNL(E852) collaboration, found a strong signal for a $I^{G}J^{PC}=0^{-}2^{+-}$ state near 1.9 GeV (labeled $h_{2}$) in the process $\pi^{-}p\rightarrow h_{2}n\rightarrow b_{1}\pi n\rightarrow 2\pi\omega\rightarrow 5\pi n$ Adams:2005tx . This could be a candidate for our isovector $2^{+-}$ state, whose mass is very close to 1.9 GeV and whose main decay mode is the $b_{1}(1235)\pi$ channel. However, these preliminary results need confirmation and should therefore be taken with caution. Searches at GlueX and in multi-pion final states in $J/\psi$ and $\chi_{c1}$ decay at BES appear to be promising for providing such confirmation.
@@
The most obvious continuation of the present work would be to compute the strong decays of the remaining states in Table 2. These hybrid states all have conventional quantum numbers ($0^{++}$$2^{++}$$1^{+-}$, $3^{+-}$) and can therefore mix with conventional mesons, making their identification more challenging. Hence, computing their strong decays might reveal some interesting features that could distinguish them from conventional mesons and thus aid in their detection. Additionally, the model presented here can be extended to include spin-dependent terms, which in turn could be used to compute the mass spectrum of hybrids.